# Unified Modeling and Structural Optimization of Multi-magnet Embedded Soft Continuum Robots for Enhanced Kinematic Performances

### v1

@@ II-B2 Magnetic Potential Energy
The proofs are deferred to Appendix A. ∎
II-C Analysis and Validation of Equilibrium Solution
[…]
TABLE I: Geometric and Material properties of the MeSCR#1
[…]
$I$ (Area moment of inertia) $0.7854$ $[\mbox{mm}^{4}]$
[…]
$M$ (Modulus of the magnetic dipole moment) $10^{-2}$ $[\mbox{A}\cdot\mbox{m}^{2}]$
We first analyze the equilibrium solution. Assuming the robot to be naturally straight $(\bar{\boldsymbol{\theta}}=\mathbf{0})$, and $\boldsymbol{\tau}(\boldsymbol{\theta})$ is ruled out. Then, reformulate (7) derives the following fixed-point equation:
$\boldsymbol{\theta}=\boldsymbol{\Lambda}^{-1}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}.$ (10)
Observe that the right-hand side of the equation represents a continuous mapping with respect to $\boldsymbol{\theta}$. We present the following theorem regarding the uniqueness of the solution.
@@ II-C Analysis and Validation of Equilibrium Solution
The inequality provides a sufficient condition to guarantee the uniqueness of a globally attractive equilibrium point for any initial condition. Notably, the upper bound is solely determined by the material properties: $E$, $I$, $L$, $M_{k}$, and is independent of the joint number $N$ (note that $N\bar{\ell}_{i}\approx L)$. The limitation on the external magnetic field strength can be explained by the critical buckling instability of the magneto-elastic material [31]. We found it intriguing that when each joint is magnetized, the resulting upper bound exhibits a striking structural resemblance to that derived in our previous work [32] on magnetic continua, differing only slightly in the coefficients. This similarity highlights the underlying consistency in the physical principles governing continuum and discreteness. Furthermore, Theorem 1 also ensures $\mathbf{S}$ to be invertible. By applying Weyl’s inequality, one can easily verify that $\lambda_{\min}(\mathbf{S})\geq\lambda_{\min}(\boldsymbol{\Lambda})-\mathcal{L}>0$.
Although the upper bound remains a conservative estimate, once the external magnetic field strength exceeds this upper limit, (10) may admit multiple equilibria pertaining to different initial conditions [21]. Therefore, selecting appropriate initial values for the iteration becomes crucial to ensuring convergence and avoiding undesirable behaviors. The solution trajectory can be described by the constructed dynamic system $\frac{\mathrm{d}\boldsymbol{\theta}}{\mathrm{d}t}=-\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}}$ with initial condition $\boldsymbol{\theta}(0)$, which can also be regarded as a gradient flow that minimizes the energy function $\mathcal{E}(\boldsymbol{\theta})$. We propose the following theorem and lemmas, which guarantee that the solution is bounded and convergent by choosing an appropriate initial condition.
Lemma 1.
For any initial condition $\boldsymbol{\theta}(0)$, the solution trajectory $\boldsymbol{\theta}(t)$ is bounded that
$\|\boldsymbol{\theta}(t)\|\leq\max\left\{\|\boldsymbol{\theta}(0)\|,\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}\right\}.$
Proof.
Consider the Lyapunov function $V(\boldsymbol{\theta})=\frac{1}{2}\|\boldsymbol{\theta}\|^{2}$, the derivative of $V(\boldsymbol{\theta})$ with respect to time is $\dot{V}(\boldsymbol{\theta})=\boldsymbol{\theta}^{\top}(-\boldsymbol{\Lambda}\boldsymbol{\theta}+\mathbf{M}(\boldsymbol{\theta})\mathbf{b})=-\boldsymbol{\theta}^{\top}\boldsymbol{\Lambda}\boldsymbol{\theta}+\boldsymbol{\theta}^{\top}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$. Using Proposition 1, the upper bound of $\dot{V}(\boldsymbol{\theta})$ is derived as $\dot{V}(\boldsymbol{\theta})\leq-\lambda_{\min}(\boldsymbol{\Lambda})\|\boldsymbol{\theta}\|^{2}+\mathcal{M}\|\boldsymbol{\theta}\|.$ Let $W(t)=\sqrt{2V(\boldsymbol{\theta})}=\|\boldsymbol{\theta}\|$, one can find out that: $\dot{W}(t)=\frac{\dot{V}(\boldsymbol{\theta})}{W(t)}\leq-\lambda_{\min}W(t)+\mathcal{M}.$ Applying Grönwall’s inequality, we obtain $W(t)\leq W(0)e^{-\lambda_{\min}(\boldsymbol{\Lambda})t}+\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}\left(1-e^{-\lambda_{\min}\left(\boldsymbol{\Lambda}\right)t}\right).$ Observed that as $t\to\infty$, the exponential term vanishes, leading to $\lim_{t\to\infty}W(t)\leq\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}.$ This completes the proof, establishing that $\|\boldsymbol{\theta}(t)\|$ remains $\|\boldsymbol{\theta}(t)\|\leq\max\left\{\|\boldsymbol{\theta}(0)\|,\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}\right\}.$
Lemma 2.
Let $\boldsymbol{\theta}^{*}$ be any equilibrium point of (10). Using Proposition 1, we have $\|\boldsymbol{\theta}^{*}\|=\|\boldsymbol{\Lambda}^{-1}\mathbf{M}(\boldsymbol{\theta}^{*})\mathbf{b}\|\leq\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}$.
Theorem 2.
The solution trajectory $(\boldsymbol{\theta}(t))$ will always converge to the largest invariant subset of the equilibrium point set if the initial condition satisfies $\|\boldsymbol{\theta}(0)\|\leq\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}$.
Proof.
Choose the Lyapunov function $V(\boldsymbol{\theta})=\mathcal{E}(\boldsymbol{\theta})$, whose derivative along the solution trajectory satisfies $\dot{V}(\boldsymbol{\theta})=-\|\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}}\|^{2}\leq 0$, with equality holding if and only if $\boldsymbol{\theta}$ is an equilibrium point. Consider the positively invariant set $\Omega=\left\{\boldsymbol{\theta}\in\mathbb{R}^{3N}\mid\|\boldsymbol{\theta}\|\leq\frac{\mathcal{M}}{\lambda_{\min}(\boldsymbol{\Lambda})}\right\}$ (Lemma 1), which contains all equilibrium points (Lemma 2). By Brouwer’s fixed-point theorem, there exists at least one equilibrium point within $\Omega$, that is, $\dot{V}(\boldsymbol{\theta})\leq 0$ holds for all $\boldsymbol{\theta}\in\Omega$. According to LaSalle’s invariance principle, every trajectory within $\Omega$ converges to the largest invariant subset where $\dot{V}(\boldsymbol{\theta})=0$. ∎
When the external magnetic field changes, requiring the recalculation of the equilibrium point, Theorem 2 provides a natural choice for the initial condition as $\boldsymbol{\theta}(0)=\mathbf{0}$ because $\forall\mathbf{b}\in\mathbb{B},\mathcal{M}\geq 0$. One can also choose the nearby solution as an initial guess if $\mathbf{b}$ continuously increases. For small perturbations on $\mathbf{b}$, the corresponding $\delta\boldsymbol{\theta}$ can be calculated by (8). Moreover, from Proposition 1 and the fact that $\boldsymbol{\theta}=\mathbf{0}$ is the unique equilibrium solution when $\mathbf{b}=\mathbf{0}$, we deduce that $\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}}$ is Lipschitz continuous with respect to $\mathbf{b}$, and the equilibrium points depends continuously on it.
[…]
Notably, using fewer joints does not introduce significant truncation errors and offers a computational speed advantage over BVP-solving models. Fig. 3e and Fig. 3f show the root-mean-square error (RMSE) of the centerline position and the distal pose error norm as the joint count increases from $3$ to $100$, benchmarked against $200$ joints. Results demonstrate that RMSE decreases proportionally to $1/N^{2}$, with an error bound established in our previous work [34]. Using $7$ joints achieves an RMSE below the acceptable $10^{-4}$ threshold. Additionally, the distal pose error norm declines rapidly as $N$ grows, confirming that fewer joints can accurately compute the distal pose under external magnetic fields. This analysis leads to a key insight: the performance index of the actuation Jacobian is largely insensitive to $N$ and can be understood as an inherent property of the robot in a specific configuration.
III Enhancing Kinematic Performances
This section focuses on the structural optimization of MeSCRs to enhance kinematic performances, including manipulability and dexterity. Prior to the optimization, we perform some kinematic analysis of the actuation Jacobian, uncovering how different magnet configuration shape the robot’s controllable DoF achievable workspace.
III-A Kinematics Analysis
@@
We define the controllable degrees of freedom (DoF) as the independently actuated direction of the task-space twist, which corresponds to the row rank of the actuation Jacobian ($\mathbf{J}_{\mathbf{b}}$). The following theorem reveals the relation between it and the configuration of embedded magnets:
@@
Recall that the actuation Jacobian $\mathbf{J}_{\mathbf{b}}=\mathbf{J}_{\boldsymbol{\theta}}\mathbf{S}^{-1}\mathbf{M}(\boldsymbol{\theta})$, the post-multiplied matrix $(\mathbf{M}(\boldsymbol{\theta}))$ can be rewritten as
$\mathbf{M}(\boldsymbol{\theta})$ $=\mathbf{D}_{k_{N_{m}-1}}^{\top}\left(\begin{bmatrix}\begin{matrix}\mathbf{1}_{k_{0}}\\ \mathbf{0}_{N-k_{0}}\end{matrix}\cdots\begin{matrix}\mathbf{1}_{k_{N_{m}-1}}\\ \mathbf{0}_{N-k_{N_{m}-1}}\end{matrix}\end{bmatrix}\otimes\mathbf{I}_{3\times 3}\right)$ (12)
×$\times\operatorname{blkdiag}\left\{\left[\mathbf{m}_{k_{0}}\right]_{\times},\dots,\left[\mathbf{m}_{k_{N_{m}-1}}\right]_{\times}\right\}$
in which the block diagonal matrix $\mathbf{D}_{k_{N_{m}-1}}$ is provided in (24) and $\mathbf{1}_{k}$ denotes the all-ones vector of $k$ dimensions. To determine the rank of $\mathbf{M}(\boldsymbol{\theta})$, one starts by noting that
$\operatorname{rk}\left(\begin{bmatrix}\begin{matrix}\mathbf{1}_{k_{0}}\\ \mathbf{0}_{N-k_{0}}\end{matrix}\cdots\begin{matrix}\mathbf{1}_{k_{N_{m}-1}}\\ \mathbf{0}_{N-k_{N_{m}-1}}\end{matrix}\end{bmatrix}\otimes\mathbf{I}_{3}\right)=3N_{m}.$
Applying Sylvester’s rk inequality, we obtain:
$\operatorname{rk}\left(\mathbf{M}(\boldsymbol{\theta})\right)$ $\leq\min\!\left\{\sum_{i=0}^{N_{m}-1}\!\operatorname{rk}\!\left(\mathbf{R}_{0}^{k_{i}}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{k_{i}}}\right),\sum_{i=0}^{N_{m}-1}\!\operatorname{rk}\!\left(\left[\mathbf{m}_{k_{i}}\right]_{\times}\right)\right\}$ (13)
$=2N_{m},$
and
$\operatorname{rk}\!\left(\mathbf{M}(\boldsymbol{\theta})\right)$ $\geq\sum_{i=0}^{N_{m}-1}\!\left(\operatorname{rk}\!\left(\mathbf{R}_{0}^{k_{i}}\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{k_{i}}}\right)+\!\operatorname{rk}\!\left(\left[\mathbf{m}_{k_{i}}\right]_{\times}\right)\right)-3N_{m}$ (14)
$=2N_{m}.$
Here we used the facts that $\mathbf{J}_{\mathbf{R}_{r}}^{\boldsymbol{\theta}_{k_{i}}}$ and $\mathbf{R}_{0}^{k_{i}}$ are both invertible and the rank of a skew-symmetric matrix remains $2$. Combined (13) with (14) leads to the following result:
$\operatorname{rk}\left(\mathbf{M}(\boldsymbol{\theta})\right)=2N_{m}.$
Consequently, an important conclusion is derived as $\operatorname{rk}\left(\mathbf{J}_{\mathbf{b}}\right)\leq\min\left\{6,\operatorname{rk}\left(\mathbf{M}(\boldsymbol{\theta})\right)\right\}=\min\left\{6,2N_{m}\right\}$. ∎
We proceed with a brief discussion on MeSCR’s structural configuration with different $N_{m}$:
Figure 4: Deformation of the MeSCR with different numbers of embedded magnets $(N_{m}=1,2,3,4)$, corresponding to subfigures a.), b.), c.), and d.). As $N_{m}$ increases, the MeSCR is characterized as underactuated for $N_{m}=1,2$, fully-actuated for $N_{m}=3$, and redundant for $N_{m}=4$.
  • $N_{m}=1$ [2, 3, 4] (see Fig. 4a.): The simplest yet most used configuration. The single magnet is typically embedded at the distal end to maximize torque transmission. From Theorem 3, we show that at most $2$ DoF of the task-space twist is controllable, which is consistent with the conclusion in [18].
  • $N_{m}=2$[35, 20, 19, 36] (see Fig. 4b.): Multimodality properties emerged since the orientation of embedded magnets is designable. However, at most $4$ DoF are controllable, which still characterizes the system as underactuated.
  • $N_{m}=3$ [37, 13, 38] (see Fig. 4c.): The structural configuration of the embedded magnets is also optimizable. The maximum controllable DoF is $6$, indicating that the MeSCR is fully-actuated.
  • $N_{m}\geq 4$[39, 40, 41] (see Fig. 4d.): Adopted by many commerical products. The column rank of $\mathbf{J}_{\mathbf{b}}$ exceeds its row rank, indicating that the MeSCR is a redundant system and the number of redundant DoF is $62N_{m}-$ 6.
The maximum controllable DoF will also be limited by the actuating field. As $N_{m}$ increases, achieving fully-actuated or redundant control requires each embedded magnet to be paired with an independent actuating field. However, generating independently controllable actuating fields at arbitrary positions within the workspace is highly challenging. While some advanced equipment may be capable of this [42], it is not feasible for most conventional systems. For example, when the actuating field is spatially uniform within the workspace, which can be expressed as $\mathbf{b}=\mathbf{U}_{u}\mathbf{b}_{u}$ with $\mathbf{U}_{u}=\mathbf{1}_{N_{m}}\otimes\mathbf{I}_{3}$ and $\mathbf{b}_{u}\in\mathbb{R}^{3}$. In this case, the maximum controllable DoF will not exceed $3$. To reduce the number of design variables, we assume all embedded magnets have uniform material and geometric properties, which also facilitates the manufacturing process. Consequently, the orientation and position of the embedded magnets become the primary variables to be optimized.
In existing commercial products, embedded magnets are generally placed along the catheter’s axial direction. For $N_{m}$ magnets, this arrangement leads to $2^{N_{m}}$ possible permutations. However, if the actuating magnetic field of each magnet is independently controllable, the directions of the magnetic moments become redundant. This redundancy arises because of the relationship: [±$[\pm\mathbf{m}_{k}]_{\times}\mathbf{b}_{k}=[\mathbf{m}_{k}]_{\times}(\pm\mathbf{b}_{k})$, indicating that the effect of flipping a magnetic moment’s direction can always be compensated by reversing the corresponding magnetic field direction. Based on this reasoning, we assume that the magnetic field can only undergo an overall direction reversal. For systems driven by electromagnets, this can be easily achieved by altering the current direction.
Aligning the magnetic moments along the axial direction significantly reduces the optimization search space but imposes a limitation on the motion of the MeSCR. Specifically, this arrangement restricts the MeSCR to material-twist-free due to the decoupling of twisting and bending, as ensured by the isotropic and axially symmetric properties of its soft matrix.
@@
The MeSCR is material-twist-free ($\boldsymbol{\theta}_{i}^{\top}\bar{\mathbf{t}}_{i}=0$) for any actuating magnetic field, provided that the embedded magnets’ magnetic moments are aligned axially, i.e., $\bar{\mathbf{m}}_{k}=\pm M_{k}\bar{\mathbf{t}}$, and the magnetic field strength satisfies:
$\max_{k\in\mathbb{K}}B_{k}\leq\frac{\sqrt{6}\lambda_{\min}(\boldsymbol{\Lambda})}{\sum_{k\in\mathbb{K}}M_{k}}.$ (15)
@@
The proof is detailed in the supplementary material. ∎
Noticing that $\frac{\sqrt{6}\lambda_{\min}(\boldsymbol{\Lambda})}{\sum_{k\in\mathbb{K}}M_{k}}\propto N^{2}$, as $N\to\infty$, Theorem 4 holds for any actuating magnetic field. This upper bound characterizes the ability of the discrete model with $N$ joints to approximate the continuum, as discussed in Section II-C. We found an alternative form of the equilibrium equation (10) from the underlying proof:
@@
For any initial condition satisfying the assumptions of Theorem 2, the associated equilibrium state is guaranteed to exhibit zero torsional components. The configuration space (1) thus can be degenerated to a simplified model as presented in [16]. Moreover, the joint stiffness matrix reduces to $\boldsymbol{\Lambda}_{i}=\Lambda_{i}\mathbf{I}_{3}$ with $\Lambda_{i}=\frac{E_{i}I_{i}}{\ell_{i}}$, and we denote $\lambda_{\min}(\boldsymbol{\Lambda})=\Lambda_{\min}$. We emphasize its distinction from geometric-twist-free, which requires $\boldsymbol{\theta}_{i}^{\top}\mathbf{t}_{i}=0$. It should be known that geometric torsion may still exist under the material-twist-free condition. However, if the external magnetic field is coplanar, the following corollary holds:
@@
Define the plane $\Pi=\operatorname{span}\{\bar{\mathbf{t}},\mathbf{v}\}\subset\mathbb{R}^{3}$, where $\mathbf{v}\in\mathbb{R}^{3}$ is linearly independent of $\bar{\mathbf{t}}$. Suppose $\forall k\in\mathbb{K}$, $\bar{\mathbf{m}}_{k}=\pm M_{k}\bar{\mathbf{t}}$ and $\mathbf{b}_{k}\in\Pi$. Then there exists an equilibrium configuration $\boldsymbol{\theta}^{*}\in\mathbb{R}^{3N}$ such that the centerline of the MeSCR lies entirely within the plane $\Pi$.
@@
[…]
Specially, if the initial value is chosen as $\boldsymbol{\theta}(0)=\mathbf{0}$, then the gradient-based iteration evolves within the subspace $\operatorname{span}\{\mathbf{n}\}$. By Theorem 2, the optimization trajectory will converge to a planar equilibrium configuration satisfying $\boldsymbol{\theta}_{i}=\theta_{i}\mathbf{n}$. We can further derive a closed-form expression under the small-angle premise, wherein the cumulative rotation matrix admits the first-order expansion: $\mathbf{R}\left(\sum_{i=0}^{k}\theta_{i}\mathbf{n}\right)\approx\mathbf{I}_{3}+\sum_{i=0}^{k}\theta_{i}[\mathbf{n}]_{\times}$. Substituting this into the total potential energy yields
@@
$\theta_{i}=\frac{1}{\Lambda_{i}}\sum_{k\in\mathbb{K},k\geq i}\bar{\mathbf{m}}_{k}^{\top}[\mathbf{b}_{k}]_{\times}\mathbf{n}.$ (17)
@@
III-B Optimization Methodology
Figure 5: Comparison of the half-planar normalized workspace of MeSCRs with equidistant placement of three embedded magnets.

### v2

@@ IV-B Properties of Magnetic Actuation under Uniform Fields
The proof is deferred to supplementary materials. ∎
Assuming the robot to be naturally straight $(\bar{\boldsymbol{\theta}}=\mathbf{0})$, and $\boldsymbol{\tau}(\boldsymbol{\theta})$ is ruled out, (16) further simplifies to the following fixed-point equation:
$\boldsymbol{\theta}=\boldsymbol{\Lambda}^{-1}\mathbf{M}(\boldsymbol{\theta})\mathbf{b}$ (17)
Under this formulation, the existence and uniqueness of the equilibrium solution can be guaranteed by ensuring that the mapping is a contraction in the configuration space:
@@
The inequality provides a sufficient condition to guarantee the uniqueness of a globally attractive equilibrium point for any initial condition. Notably, the upper bound is solely determined by the material properties and is independent of the joint number $N$ (note that $N\bar{\ell}_{i}\approx L$). The limitation on the external magnetic field strength can be explained by the critical buckling instability of the magneto-elastic material [27]. We found it intriguing that when each joint is magnetized, the resulting upper bound exhibits a striking structural resemblance to that derived in previous work [28] on magnetic continua, differing only slightly in the coefficients. Furthermore, Theorem 1 also ensures $\mathbf{S}$ to be invertible. By applying Weyl’s inequality, one can easily verify that $\lambda_{\min}(\mathbf{S})\geq\lambda_{\min}(\boldsymbol{\Lambda})-\mathcal{L}>0$.
@@
We subsequently define the controllable degrees of freedom (DoF) as the independently actuated direction of the task-space twist, which corresponds to the row rank of the actuation Jacobian $(\mathbf{J}_{\mathbf{b}}=\mathbf{J}_{\boldsymbol{\theta}}\frac{\partial\boldsymbol{\theta}}{\partial\mathbf{b}})$. The following theorem reveals the relation between it and the configuration of embedded magnets:
@@
It is deferred to the supplementary materials. ∎
The controllable DoF is ultimately constrained by the number and independence of the actuating fields. As $N_{m}$ increases, achieving fully actuated or redundant control would require an equal number of independently controllable fields, which is generally infeasible in practice. Under the spatially uniform condition on fields, the maximum controllable DoF does not exceed three. For simplicity and manufacturability, all embedded magnets are assumed to share identical material and geometric properties.
In most commercial designs, embedded magnets are placed along the catheter’s axial direction. For $N_{m}$ magnets, this yields $2^{N_{m}}$ orientation permutations. However, because $\pm\mathbf{m}_{k}]_{\times}\mathbf{b}_{k}=[\mathbf{m}_{k}]_{\times}(\pm\mathbf{b}_{k})$, flipping a magnet’s polarity can be equivalently compensated by reversing the field direction. Thus, axial alignment effectively reduces the optimization space while restricting the MeSCR to a twist-free deformation mode due to the isotropic and axially symmetric nature of its soft matrix.
@@
The MeSCR is material-twist-free ($\boldsymbol{\theta}_{i}^{\top}\bar{\mathbf{t}}_{i}=0$) for any actuating magnetic field, provided that the embedded magnets’ magnetic moments are aligned axially, i.e., $\bar{\mathbf{m}}_{k}=\pm M_{k}\bar{\mathbf{t}}$, and the magnetic field strength satisfies: $\max_{k\in\mathbb{K}}B_{k}\leq\sqrt{6}\lambda_{\min}(\boldsymbol{\Lambda})/\sum_{k\in\mathbb{K}}M_{k}.$
@@
See supplementary material for details. ∎
Noticing that $\frac{\sqrt{6}\lambda_{\min}(\boldsymbol{\Lambda})}{\sum_{k\in\mathbb{K}}M_{k}}\propto N^{2}$, as $N\to\infty$, Theorem 3 holds for any actuating magnetic field. This upper bound characterizes the ability of the discrete model with $N$ joints to approximate the continuum. We found an alternative form of the equilibrium equation (17) from the underlying proof:
@@
The configuration space (1) thus can be degenerated to a simplified model as presented in [29]. Moreover, the joint stiffness matrix reduces to $\boldsymbol{\Lambda}_{i}=\Lambda_{i}\mathbf{I}_{3}$ with $\Lambda_{i}=\frac{E_{i}I_{i}}{\ell_{i}}$, and we denote $\lambda_{\min}(\boldsymbol{\Lambda})=\Lambda_{\min}$. We emphasize its distinction from geometric-twist-free, which requires $\boldsymbol{\theta}_{i}^{\top}\mathbf{t}_{i}=0$. It should be known that geometric torsion may still exist under the material-twist-free condition. However, if the external magnetic field is coplanar, the following corollary holds:
@@
Define the plane $\Pi=\operatorname{span}\{\bar{\mathbf{t}},\mathbf{s}\}\subset\mathbb{R}^{3}$, where $\mathbf{s}\in\mathbb{R}^{3}$ is linearly independent of $\bar{\mathbf{t}}$. Suppose $\forall k\in\mathbb{K}$, $\bar{\mathbf{m}}_{k}=\pm M_{k}\bar{\mathbf{t}}$ and $\mathbf{b}_{k}\in\Pi$. Then there exists an equilibrium configuration $\boldsymbol{\theta}^{*}\in\mathbb{R}^{3N}$ such that the centerline of the MeSCR lies entirely within the plane $\Pi$.
@@
The proof is constructive. Let $\boldsymbol{\theta}_{i}^{*}=\theta_{i}^{*}\mathbf{n}$, where $\mathbf{n}=\frac{\bar{\mathbf{t}}\times\mathbf{s}}{\|\bar{\mathbf{t}}\times\mathbf{s}\|}$ is the unit vector orthogonal to $\Pi$. Since $\mathbf{b}_{k}\in\Pi$, we notice that $\mathbf{R}_{0}^{k^{\top}}(\boldsymbol{\theta}^{*})\mathbf{b}_{k}=\mathbf{R}(-\sum_{i}\theta_{i}^{*}\mathbf{n})\mathbf{b}_{k}\in\Pi$. Then $\forall k\in\mathbb{K}$, $\bar{\mathbf{m}}_{k}]_{\times}\mathbf{R}_{0}^{k^{\top}}\mathbf{b}_{k}\!\in\operatorname{span}\{\mathbf{n}\}$. Hence $\boldsymbol{\theta}_{i}^{*}=\theta_{i}^{*}\mathbf{n}$ is a solution to (19). The corresponding tangent vectors $\mathbf{t}_{i}=\mathbf{R}_{0}^{i}\bar{\mathbf{t}}_{i}\in\Pi$, thus the centerline lies entierly within the plane $\Pi$. ∎
Specially, if the initial value is chosen as $\boldsymbol{\theta}(0)=\mathbf{0}$, then the gradient-based iteration evolves within the subspace $\operatorname{span}\{\mathbf{n}\}$. The optimization trajectory converges to a planar equilibrium configuration satisfying $\boldsymbol{\theta}_{i}=\theta_{i}\mathbf{n}$. We can further derive a closed-form expression under the small-angle premise, wherein the cumulative rotation matrix admits the first-order expansion: $\mathbf{R}\left(\sum_{i=0}^{k}\theta_{i}\mathbf{n}\right)\approx\mathbf{I}_{3}+\sum_{i=0}^{k}\theta_{i}[\mathbf{n}]_{\times}$. Substituting this into the total potential energy yields
@@
$\theta_{i}=\frac{1}{\Lambda_{i}}\sum_{k\in\mathbb{K},k\geq i}\bar{\mathbf{m}}_{k}^{\top}[\mathbf{n}]_{\times}^{\top}\mathbf{b}_{k}.$ (20)
@@
IV-C Analytical Optimization under Weak Uniform Fields